### v2

@@ C.1 Auxiliary lemmas
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}+\lambda\Big{(}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\quad\in[1/4,1]\cdot\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,\lambda)(1+o(1)).$
@@
The upper bound is a direct consequence of Theorem 6.3.
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}$ $\geq\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\frac{1}{4}\bar{\psi}(1+o(1))+o(1/K_{n}).$
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\geq\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\frac{1}{4}\bar{\psi}(1+o(1)).$ (52)
@@
$\geq\frac{1}{4}\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,\lambda)(1+o(1)).$
@@
From the above proof, it is not hard to show that the worst-case MSE, up to $1+o(1)$ terms, can be formulated as
@@
$\mathbb{E}\Big{[}\Big{(}\hat{\tau}_{n}-\mathbb{E}[\hat{\tau}_{n}]\Big{)}\Big{]}=\frac{4}{n^{2}}\sum_{i,j:c(i)=c(j)}\mathrm{Cov}\Big{(}D_{i},D_{j}\Big{)}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))(\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0}))+o(1/K_{n}).$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}\Big{[}\Big{(}\hat{\tau}_{n}-\mathbb{E}[\hat{\tau}_{n}]\Big{)}\Big{]}\leq\frac{4}{n^{2}}\sum_{i,j:c(i)=c(j)}\mathrm{Cov}\Big{(}D_{i},D_{j}\Big{)}\bar{\psi}+o(1/K_{n}).$
@@
C.15 Proof of Lemma B.1
First, we note that $\tilde{\epsilon}$ does not depend on $\mathbf{d}$ and
$\mathbf{C}(\mathbf{d})\equiv\mathbf{C}\triangleq\mathrm{Cor}((\mathbf{I}-\gamma_{n}\mathbf{L})^{-1}\epsilon)=\mathbf{D}_{n}^{-1/2}(\mathbf{I}-\gamma_{n}\mathbf{L})^{-2}\mathbf{D}_{n}^{-1/2},$

### v3

@@ E.2.2 Proof of the theorem
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}+\lambda\Big{(}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\quad\in[1/4,1]\cdot\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,\lambda)(1+\mathcal{O}(\varepsilon)).$
@@
The upper bound is a direct consequence of Theorem B.3.
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}$ $\geq\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\frac{1}{4}\bar{\psi}(1+\mathcal{O}(\varepsilon))+\mathcal{O}(\varepsilon/K_{n}).$
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\geq\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\frac{1}{4}\bar{\psi}(1+\mathcal{O}(\varepsilon)).$ (67)
@@
$\geq\frac{1}{4}\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,\lambda)(1+\mathcal{O}(\varepsilon)).$
@@
From the above proof, it is not hard to show that the worst-case MSE, up to $1+\mathcal{O}(\varepsilon)$ terms, can be formulated as
@@
$\mathbb{E}\Big{[}\Big{(}\hat{\tau}_{n}-\mathbb{E}[\hat{\tau}_{n}]\Big{)}\Big{]}=\frac{4}{n^{2}}\sum_{i,j:c(i)=c(j)}\mathrm{Cov}\Big{(}D_{i},D_{j}\Big{)}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))(\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0}))+\mathcal{O}(\varepsilon/K_{n}).$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}\Big{[}\Big{(}\hat{\tau}_{n}-\mathbb{E}[\hat{\tau}_{n}]\Big{)}\Big{]}\leq\frac{4}{n^{2}}\sum_{i,j:c(i)=c(j)}\mathrm{Cov}\Big{(}D_{i},D_{j}\Big{)}\bar{\psi}+\mathcal{O}(\varepsilon/K_{n}).$
@@
E.8 Proof of Lemma B.6
First, we note that $\tilde{\nu}$ does not depend on $\mathbf{d}$ and
$\mathbf{C}(\mathbf{d})\equiv\mathbf{C}\triangleq\mathrm{Cor}((\mathbf{I}-\gamma_{n}\mathbf{L})^{-1}\nu)=\mathbf{D}_{n}^{-1/2}(\mathbf{I}-\gamma_{n}\mathbf{L})^{-2}\mathbf{D}_{n}^{-1/2},$
@@
E.9 Proof of Theorem B.7
We decompose the proof as follows. First, we characterize the bias and variance of the Bernoulli design. We then compare $\mathcal{B}_{n}$ for the cluster and Bernoulli design, assuming that $\lambda\in(0,\infty)$ is bounded away from zero and infinity as assumed in Theorem B.7.
Worst-case bias and variance of the Bernoulli and cluster design
Note that $\mathcal{C}_{\mathrm{B},n}$ satisfies Assumption 4. By Lemma D.1 and D.2, $\mathcal{C}_{\mathrm{B},n}$ has worst-case bias $\bar{\phi}_{n}b_{n}(\mathcal{C}_{\mathrm{B},n})=\bar{\phi}_{n}$, and worst-case variance $\frac{1}{n}[\bar{\psi}+o(1)]$. Thus,
and
Case $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)>\mathcal{B}_{n}^{*}(\mathcal{C}_{\mathrm{B},n},\lambda)$.
By Theorem 3.3, it remains to prove that
$\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)\geq\frac{\bar{\psi}}{K_{n}},$
and
Case $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)<\mathcal{B}_{n}^{*}(\mathcal{C}_{\mathrm{B},n},\lambda)$
Similarly, by Theorem 3.3, it remains to prove that
and